OpenStudy (anonymous):

@Loveiskey18 Checking account A charges a monthly fee of $7 and a per-check fee of $0.20, and checking account B charges a monthly fee of $8 and a per-check fee of $0.15. How many checks would a person have to write for the two accounts to cost the same? A. 20 B. 15 C. 10 D. 30

OpenStudy (anonymous):

C=7 + 0.20t C=8 + 0.15t 7+0.20t=8+0.15+ Solve for t: 7+0.20t-0.15=8+0.15t-0.15t 7t+=8 7-7+t=8+7= What is 8+7=?

OpenStudy (anonymous):

15